Flamo:

Which of the following points lies on the circle whose center is at the origin and whose radius is 5?

Flamo:

Where is the Other Radius?

Shadow:

\[(x-h)^{2} + (y-k)^{2} = r^{2}\]

Flamo:

Huh?

Shadow:

h and k are the x and y coordinates for the center of the circle. r is the radius

Shadow:

I just noticed that it says the circle is at the origin. So then we can use this formula: \[x^{2} + y^{2} = r^{2}\]

Shadow:

I assume they gave you some points.

Shadow:

1 attachment
Shadow:

The above image demonstrates why that formula works

Flamo:

o

Flamo:

The answer Choices are (-3, 4) (1, -2) And \[\sqrt{5}, \sqrt{5} \]

Flamo:

So those are the Points?

Shadow:

Yes

Flamo:

So now what?

Shadow:

\[x^2 + y^2 = r^2\] So which points satisfy \[x^2 + y^2 = 5^2\] \[x^2 + y^2 = 25\]

Flamo:

So if Its 3 and 4, 3^2 + 4^2 = 5^2

Flamo:

Then

Shadow:

Mhm

Flamo:

Woops I meant -3 not 3

Flamo:

So Its, 9 + 16 = 25

Flamo:

9 + 16 = 25

Shadow:

Yeah those are the correct points.

Flamo:

So 3, 4 and the correct Points.

Flamo:

Ye.

Flamo:

Ty Shadow!

Shadow:

No problem